# Gravitational wave propagation in hybrid metric-Palatini gravity

### v1

@@ Abstract
In this work we analyze the propagation properties of gravitational waves in the hybrid metric-Palatini gravity theory. We introduce the scalar-tensor representation of the theory to make explicit the scalar degrees of freedom of the theory and obtain their equations of motion in a form decoupled from the metric tensor. Then, we introduce linear perturbations for the metric tensor and for the two scalar fields and obtain the propagation equations for these three quantities. We analyzed the theory both at non-linear and at linear level through the Newman-Penrose formalism so to find the polarization states. We show that the tensor modes propagate at the speed of light and feature the usual $+$– and $\times$–polarization modes also present in General Relativity (GR), plus two additional polarization modes: a longitudinal mode and a breathing mode, described by the same additional degree of freedom. On the other hand, the theory features two additional scalar modes not present in GR. These modes are massive and, thus, propagate with a speed smaller than the speed of light in general. The masses of the scalar modes depend solely on the interaction potential between the two scalar fields in the theory, which suggests that one can always fine-tune the potential to make the scalar modes massless and reduce their propagation speed to the speed of light. Given the possibility of fine-tuning the theory to match the observational predictions of GR and in the absence of any measured deviations, these features potentially render the hybrid metric-Palatini theory unfalsifiable in the context of gravitational wave propagation.
@@ I Introduction
A combined version of both the metric and the Palatini formulations of $f\left(R\right)$ gravity, known as hybrid metric-Palatini gravity Capozziello:2015lza , successfully addresses the problem of unifying the late-time cosmic acceleration phase with the weak-field solar system dynamics Harko:2011nh ; Capozziello:2013uya . Later on, the theory was proven useful to overcome a wide variety of limitations of $f\left(R\right)$ gravity in the fields of astrophysics and cosmology Capozziello:2012ny ; Capozziello:2012qt ; Koivisto:2013kwa ; Capozziello:2012hr ; Capozziello:2013yha ; Carloni:2015bua ; Edery:2019txq ; Harko:2020ibn ; Danila:2016lqx ; Danila:2018xya ; Bronnikov:2019ugl ; Leanizbarrutia:2017xyd ; Bekov:2020dww ; Chen:2020evr ; Dyadina:2023exu . A generalization of the theory to allow for an arbitrary dependence of the action in the Ricci scalars of both the metric and the Palatini formalism was also developed Tamanini:2013ltp , leading to an even broader range of applications in these fields Rosa:2018jwp ; Rosa:2017jld ; Rosa:2020uoi ; Rosa:2019ejh ; Borowiec:2020lfx ; Rosa:2021yym ; Rosa:2020uli ; Bombacigno:2019did ; Rosa:2021mln ; Sa:2020qfd ; Rosa:2021lhc ; Rosa:2021ish ; daSilva:2021dsq ; Bronnikov:2021tie ; Golsanamlou:2023wiz ; Rosa:2024pzo . In this work, we aim to extend this literature to the field of gravitational waves (GWs).
@@
The observation of GWs marked the beginning of a new era in fundamental physics research LIGOScientific:2018mvr ; KAGRA:2023pio ; Barack:2018yly ; Bailes:2021 , providing remarkable insights into the nature of compact objects and binary systems Cardoso:2017cqb ; Berti:2019wnn ; LIGOScientific:2018jsj , tests of gravity in the strong-field regime LIGOScientific:2016lio ; Yunes:2016jcc ; Berti:2018cxi ; Berti:2018vdi , and even cosmological measurements LIGOScientific:2017adf ; DES:2019ccw . Naturally, the propagation of gravitational waves was also analyzed in the context of modified theories of gravity Bernard:2022noq ; Dalang:2020eaj ; Capozziello:2006ra ; Katsuragawa:2019uto ; Gong:2017bru ; Abedi:2017jqx ; Capozziello:2008 . Interestingly however, only a few works have addressed this topic in hybrid metric-Palatini gravity Dyadina:2022elo ; Dyadina:2024jjo ; Kausar:2018ipo . The results presented in these publications are shallow: the polarization modes are obtained from a purely perturbative approach, and none of the works take into consideration the generalized version of the theory. This work aims to suppress these gaps by extending the analysis to the generalized version of the theory and including an analysis of the polarization modes based on the Newman-Penrose (NP) formalism NP .
@@
The NP formalism is based on a tetrad basis, and it is useful to analyze black hole and gravitational wave solutions. For metric theories of gravity that allow for plane null gravitational wave solutions, one can resort to the Eardley-Lee-Lightman (ELL) formalism which simplifies the former and allows for the classification of the spacetime solutions ELL1 ; ELL2 , based on which one can analyze all six polarization states a metric theory can exhibit. This has been applied to modified gravity theories Alves:2009 ; Gomes:2018 . However, when a propagating mode is massive, the ELL formalism cannot be applied directly, and one has to invert its definitions to correctly find the polarization states Polarisations . A notable exception arises if the mass of the polarization mode is small, for which one obtains nearly null plane gravitational waves and the ELL formalism is argued to be approximately applicable Capozziello:2020 . Moreover, there is a distinction between polarization modes that affect a ring of test particles when a gravitational wave passes through and the independent radiative degrees of freedom a theory may exhibit; hence, for instance, $f(R)$ theories possess three radiative degrees of freedom which can excite up to four polarization states Alves:2024 .
@@
This manuscript is organized as follows. In Sec. II we introduce the generalized version of the hybrid metric-Palatini gravity in both the geometrical and scalar-tensor representations and obtain their respective decoupled equations of motion; in Sec. III we introduce linear perturbations of the metric and scalar fields to obtain the propagation equations and analyze the polarization modes of plane gravitational waves; in Sec. IV we introduce the Newman-Penrose formalism to determine the polarization states and degrees of freedom of gravitational waves when massive modes are present; in Sec. V we analyze the results obtained from both methods above and their implications for the theory from an observational point of view; and in Sec. VI we discuss our results and trace our conclusions.
@@ II Theory and equations
II.1 Hybrid metric-Palatini gravity
The action $S$ that describes the generalized hybrid metric-Palatini theory of gravity (GHMPG) is given by
@@ II.2 Scalar-tensor representation
$S$ $=$ $\frac{1}{2\kappa^{2}}\int_{\Omega}\sqrt{-g}\left[\left(\varphi-\psi\right)R-\right.$ (3)
@@
$V\left(\varphi,\psi\right)=-f\left(R,\mathcal{R}\right)+\varphi\frac{\partial f}{\partial R}-\psi\frac{\partial f}{\partial\mathcal{R}}.$ (5)
We note that the scalar-tensor representation is only well-defined if Eqs. (4) are invertible, i.e., if it is possible to write $R=R\left(\varphi,\psi\right)$ and $\mathcal{R}=\mathcal{R}\left(\varphi,\psi\right)$.
The modified field equations in the scalar-tensor representation can be obtained through the variation of Eq. (3) with respect to the metric $g_{\mu\nu}$, which take the form
@@ III.2 Tensor modes
$2\nabla^{\alpha}\nabla_{(\mu}\chi_{\nu)\alpha}-\eta_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\chi^{\alpha\beta}-\Box\chi_{\mu\nu}=0$ (26)
@@
where $e^{\pm}_{\mu\nu}$ represent the amplitudes of the outgoing ($+$) and ingoing (--) waves, and $k^{\mu}$ is the wave vector. The wave vector and the amplitudes satisfy the orthogonality conditions
@@
The metric perturbation in Eq. (32) presents three polarization modes described by the metric perturbation components $\delta g_{yy}$, $\delta g_{zz}$, and $\delta g_{yz}$. The metric component $\delta g_{yz}$ characterizes the usual $\times$-polarization also present in GR. However, the components $\delta g_{yy}$ and $\delta g_{zz}$ cease to be symmetric, and thus they contribute with both a transversal and a longitudinal polarizations. Via the imposition of the GR condition $\delta g=0$, from which one obtains $\delta g_{yy}=-\delta g_{zz}$, the additional longitudinal polarization mode is eliminated and the additional transversal polarization mode reduces to the usual $+$-polarization in GR.
@@ III.3 Scalar modes
where $\Phi^{\pm}$ and $\Psi^{\pm}$ represent the amplitudes of the outgoing ($+$) and ingoing (--) waves, and $p^{\mu}$ and $q^{\mu}$ are the wave vectors. Due to the fact that the scalar modes are massive, the normalization conditions for the wave vectors in this case become
@@
$M_{\Phi}=\frac{1}{3}\left(A-\sqrt{B}\right),$ (42)
$M_{\Psi}=\frac{1}{3}\left(A+\sqrt{B}\right),$
@@ IV.1 The Newman-Penrose formalism
$R_{\rho\sigma\mu\nu}$ $=$ $C_{\rho\sigma\mu\nu}+\frac{1}{2}\left(g_{\sigma\nu}R_{\rho\mu}+g_{\rho\mu}R_{\sigma\nu}-g_{\sigma\mu}R_{\rho\nu}-\right.$ (47)
@@
Eardley, Lee, and Lightman have developed a formalism for classifying the gravitational wave solution for metric theories allowing for plane null gravitational wave solutions, that is, with massless modes, in which the NP quantities reduce to only six real independent components, namely the set $\{\Psi_{2},~{}\Psi_{3},~{}\Psi_{4},~{}\Phi_{22}\}$ (the other nonvanishing components can be shown to be linear combinations of the elements of this set). The associated symmetry group is a subgroup of the Lorentz group known as the “little group” $E(2)$ that leaves the wave-vector invariant. Under the action of the rotation group, the values of the helicity for each of the components of the set are, respectively, $\{0,~{}\pm 1,\pm 2,0\}$. Hence, $\Psi_{2}$ represents a longitudinal mode, $\Psi_{3}$ can be split into the real and purely imaginary parts responsible for the vectorial $y-$ and $z-$ modes, $\Psi_{4}$ can also be split as the former leading to the tensor $+$ and $\times$ polarizations, and $\Phi_{22}$ represents the scalar breathing mode ELL1 ; ELL2 .
For gravity theories with massive gravitational wave polarization modes, the formalism outlined above cannot be straightforwardly applied. As we have found in Sec. III through the linearization method, the hybrid metric-Palatini theory features two massive scalar modes that can be decoupled. Thus, we shall use the method of inverting the NP quantities as functions of the polarization states as done in Ref. Polarisations .
@@ IV.2 Massive Propagating Modes
where $p_{1},\dots,p_{6}$ are the longitudinal, vectorial y-, vectorial z-, $+$, $\times$, and breathing polarization modes, respectively.
@@
Equations (59) represent the most general description of gravitational wave polarization modes in the hybrid metric-Palatini gravity theory, since the assumption of propagation in the $x$ direction is arbitrary. To proceed with the analysis of these modes, it is necessary to break generality either imposing explicit forms for the quantities appearing in the equations, e.g. choosing a particular form of the theory, or to impose additional constraints on these quantities through a comparison with observational data.
@@ V Constraints and implications
For the particular case of plane gravitational waves propagating along the $x$ direction, which we consider to be an acceptable approximation given the large distance to the sources of the gravitational wave signals detected, the displacement is constant along each plane orthogonal to the propagation direction. Hence we expect that $g_{xy}=0=g_{xz}$, a result that is in agreement with the linearized case (see Eq. (32)), and therefore no vector polarization modes arise in the full non-linear regime of the theory, i.e. $p_{2}^{(y)}=0=p_{3}^{(z)}$.
Additional simplifications to the results presented in Eqs. (59) can be obtained by taking into account the constrains on the scalar fields obtained from a cosmological analysis of the theory. Indeed, this gravity model has been thoroughly studied in a cosmological context through the dynamical systems approach Rosa:2024pzo , and recent results show that the fields $\varphi$ and $\psi$ are exponentially suppressed during the matter-dominated epoch, reaching approximately constant values at late times, i.e. $\varphi,\psi\approx const$. Furthermore, the solar system dynamics constrain the scalar fields to $\varphi-\psi\approx 1$. As such, the potential $V$ is responsible for late-time cosmic acceleration and plays the role of a cosmological constant at late times, whose value is constrained to $V\approx 10^{-122}/m_{P}^{2}$, with $m_{P}$ the Planck mass $m_{P}=\sqrt{\hbar G/c^{3}}$.
Taking the cosmological results into account at non-linear level, we expect that at present times, since $\partial_{\mu}\varphi\approx 0$, $\partial_{\mu}\psi\approx 0$, and $V\sim 10^{-122}$, and thus $\Xi\approx 0$, the theory may exhibit highly-suppressed scalar modes, $p_{1}^{(l)}$, $p_{6}^{(b)}$, and tensor polarizations which are essentially dominated by the Riemann tensor components. Even so, it is remarkable that the $+$ polarization is not necessarily identical to the one in GR, as it is not required that $R_{xyxy}=-R_{xzxz}$, and hence we there is a possible ”smoking gun” for distinguishing between GHMPG and GR. However, it is important to note that since GR exists as a particular solution of GHMPG, in the absence of potentially measurable differences between $R_{xyxy}$ and $-R_{xzxz}$, this analysis renders the hybrid metric-Palatini gravity unfalsifiable from the point of view of gravitational wave propagation in the plane gravitational wave approximation.
@@
$p_{4}^{(+)}$ =±$=\pm\frac{1}{2}\omega^{2}\left(\delta g_{yy}-\delta g_{zz}\right)~{},$ (62)
$p_{5}^{(\times)}$ =±$=\pm\frac{1}{2}\omega^{2}\delta g_{yz}~{},$
$p_{6}^{(b)}$ $=-\frac{1}{2}\left(\delta g_{yy,tt}+\delta g_{zz,tt}\right)~{},$
In fact, one could fully estimate the longitudinal mode in the linear limit, by using Eqs. (38) and (40) in order to obtain

### v2

# Gravitational wave propagation in generalized hybrid metric-Palatini gravity

@@ Abstract
In this work we analyze the propagation properties of gravitational waves in the generalized hybrid metric-Palatini gravity theory. We introduce the scalar-tensor representation of the theory to make explicit the scalar degrees of freedom of the theory and obtain their equations of motion in a form decoupled from the metric tensor. Then, we introduce linear perturbations for the metric tensor and for the two scalar fields and obtain the propagation equations for these three quantities. We analyzed the theory both at non-linear and at linear level through the Newman-Penrose formalism so to find the polarization states. We show that the tensor modes propagate at the speed of light and feature the usual ++– and $\times$–polarization modes also present in General Relativity (GR), plus two additional polarization modes: a longitudinal mode and a breathing mode, described by the same additional degree of freedom. On the other hand, the theory features two additional scalar modes not present in GR. These modes are massive and, thus, propagate with a speed smaller than the speed of light in general. The masses of the scalar modes depend solely on the interaction potential between the two scalar fields in the theory, which suggests that one can always fine-tune the potential to make the scalar modes massless and reduce their propagation speed to the speed of light. Given the possibility of fine-tuning the theory to match the observational predictions of GR and in the absence of any measured deviations, these features potentially render the generalized hybrid metric-Palatini theory unfalsifiable in the context of gravitational wave propagation.
@@ I Introduction
A combined version of both the metric and the Palatini formulations of $f\left(R\right)$ gravity, known as hybrid metric-Palatini gravity Capozziello:2015lza , successfully addresses the problem of unifying the late-time cosmic acceleration phase with the weak-field solar system dynamics Harko:2011nh ; Capozziello:2013uya . Later on, the theory was proven useful to overcome a wide variety of limitations of $f\left(R\right)$ gravity in the fields of astrophysics and cosmology Capozziello:2012ny ; Capozziello:2012qt ; Koivisto:2013kwa ; Capozziello:2012hr ; Capozziello:2013yha ; Carloni:2015bua ; Edery:2019txq ; Harko:2020ibn ; Danila:2016lqx ; Danila:2018xya ; Bronnikov:2019ugl ; Leanizbarrutia:2017xyd ; Bekov:2020dww ; Chen:2020evr ; Dyadina:2023exu . A generalization of the theory to allow for an arbitrary dependence of the action in the Ricci scalars of both the metric and the Palatini formalism was also developed Tamanini:2013ltp (which is known as generalized hybrid metric-Palatini gravity, or GHMPG), leading to an even broader range of applications in these fields Rosa:2018jwp ; Rosa:2017jld ; Rosa:2020uoi ; Rosa:2019ejh ; Borowiec:2020lfx ; Rosa:2021yym ; Rosa:2020uli ; Bombacigno:2019did ; Rosa:2021mln ; Sa:2020qfd ; Rosa:2021lhc ; Rosa:2021ish ; daSilva:2021dsq ; Bronnikov:2021tie ; Golsanamlou:2023wiz ; Rosa:2024pzo ; Bombacigno:2024lud . In this work, we aim to extend this literature to the field of gravitational waves (GWs).
@@
The observation of GWs marked the beginning of a new era in fundamental physics research LIGOScientific:2018mvr ; KAGRA:2023pio ; Barack:2018yly ; Bailes:2021 , providing remarkable insights into the nature of compact objects and binary systems Cardoso:2017cqb ; Berti:2019wnn ; LIGOScientific:2018jsj , tests of gravity in the strong-field regime LIGOScientific:2016lio ; Yunes:2016jcc ; Berti:2018cxi ; Berti:2018vdi , and even cosmological measurements LIGOScientific:2017adf ; DES:2019ccw . Naturally, the propagation of gravitational waves was also analyzed in the context of modified theories of gravity Bernard:2022noq ; Dalang:2020eaj ; Capozziello:2006ra ; Katsuragawa:2019uto ; Gong:2017bru ; Abedi:2017jqx ; Capozziello:2008 ; Lopez:2025gfu ; Casado-Turrion:2024esi . Interestingly however, only a few works have addressed this topic in hybrid metric-Palatini gravity Dyadina:2022elo ; Dyadina:2024jjo ; Kausar:2018ipo . The results presented in these publications are shallow: the polarization modes are obtained from a purely perturbative approach, and none of the works take into consideration the generalized version of the theory. This work aims to suppress these gaps by extending the analysis to the generalized version of the theory and including an analysis of the polarization modes based on the Newman-Penrose (NP) formalism NP .
@@
The NP formalism is based on a tetrad basis, and it is useful to analyze black hole and gravitational wave solutions. For metric theories of gravity that allow for plane null gravitational wave solutions, one can resort to the Eardley-Lee-Lightman (ELL) formalism which simplifies the former and allows for the classification of the spacetime solutions ELL1 ; ELL2 , based on which one can analyze all six polarization states a metric theory can exhibit. In fact, this formalism has been applied to modified gravity theories Alves:2009 ; Gomes:2018 . However, when a propagating mode is massive, the ELL formalism cannot be applied directly, and one has to invert its definitions to correctly find the polarization states Polarisations . A notable exception arises if the mass of the polarization mode is small, for which one obtains nearly null plane gravitational waves and the ELL formalism is argued to be approximately applicable Capozziello:2020 . Moreover, there is a distinction between polarization modes that affect a ring of test particles when a gravitational wave passes through and the independent radiative degrees of freedom a theory may exhibit; hence, for instance, $f(R)$ theories possess three radiative degrees of freedom which can excite up to four polarization states Alves:2024 .
@@
This manuscript is organized as follows. In Sec. II, we introduce the GHMPG theory in both the geometrical and scalar-tensor representations and obtain their respective decoupled equations of motion; in Sec. III we introduce linear perturbations of the metric and scalar fields to obtain the propagation equations and analyze the polarization modes of plane gravitational waves; in Sec. IV we introduce the Newman-Penrose formalism to determine the polarization states and degrees of freedom of gravitational waves when massive modes are present; in Sec. V we analyze the results obtained from both methods above and their implications for the theory from an observational point of view; and in Sec. VI we discuss our results and trace our conclusions.
@@ II Theory and equations
II.1 Generalized hybrid metric-Palatini gravity
The action $S$ that describes the GHMPG theory is given by
@@ II.2 Scalar-tensor representation
[…]
@@
$V\left(\varphi,\psi\right)=-f\left(R,\mathcal{R}\right)+\varphi R-\psi\mathcal{R}$ (5)
We note that, for the transformation from the geometrical representation to the scalar-tensor representation to be well-defined, it is required that the determinant of the Hessian matrix of $f\left(R,\mathcal{R}\right)$ is non-zero. This is equivalent to requiring that Eqs. (4) are invertible. Indeed, these equations imply that the scalar fields $\varphi$ and $\psi$ can be written as functions of $R$ and $\mathcal{R}$ as $\varphi\left(R,\mathcal{R}\right)$ and $\psi\left(R,\mathcal{R}\right)$, and the transformation is well-defined if it is possible to write $R$ and $\mathcal{R}$ as functions of $\varphi$ and $\psi$ as $R\left(\varphi,\psi\right)$ and $\mathcal{R}\left(\varphi,\psi\right)$. This invertibility guarantees the uniqueness of the transformation (see Ref. Rosa:2017jld ).
The modified field equations in the scalar-tensor representation can be obtained through the variation of Eq. (3) with respect to the metric $g_{\mu\nu}$, which take the form (see e.g. Rosa:2018jwp )
@@ III.2 Tensor modes
$2\nabla^{\alpha}\nabla_{(\mu}\chi_{\nu)\alpha}-\eta_{\mu\nu}\nabla_{\alpha}\nabla_{\beta}\chi^{\alpha\beta}-\Box\chi_{\mu\nu}=0.$ (26)
@@
where $e^{\pm}_{\mu\nu}$ represent the amplitudes of the outgoing (++) and ingoing (-) waves, and $k^{\mu}$ is the wave vector. The wave vector and the amplitudes satisfy the orthogonality conditions
@@
The metric perturbation in Eq. (32) presents three degrees of freedom associated with the metric perturbation components $\delta g_{yy}$, $\delta g_{zz}$, and $\delta g_{yz}$. The metric perturbation component $\delta g_{yz}$ characterizes the usual $\times$-polarization also present in GR. However, the components $\delta g_{yy}$ and $\delta g_{zz}$ cease to be symmetric, and thus they contribute with both a transversal and a longitudinal polarizations. At this point, one can verify that the GHMPG encompasses GR as a particular solution. Indeed, when one uses the usual gauge condition in GR, $\nabla^{\mu}h_{\mu\nu}\equiv\nabla^{\mu}\left(\delta g_{\mu\nu}-\frac{1}{2}\eta_{\mu\nu}\delta g\right)=0$ (which is equivalent to $\nabla^{\mu}\chi_{\mu\nu}=0$ in our case), the analogous of Eq. (28) has a factor of $\frac{1}{2}$ instead of $\frac{1}{4}$ (the latter is the result of the uncoupling of the extra scalar degree of freedom of GHMPG from the tensor modes, a characteristic that can arise in modified gravity theories Gomes:2018 ). Therefore, in such a gauge, the metric perturbation in GR satisfies the condition $\delta g=0$ (which together with the previous transverse condition is the well-known traceless-transverse gauge, or TT-gauge). Then, if one imposes this condition on the metric perturbation given in Eq. (32), $\delta g_{tt}$ and $\delta g_{xx}$ cancel out, forcing $\delta g_{yy}=-\delta g_{zz}$. The additional longitudinal polarization mode is thus removed, and the additional transversal polarization mode reduces to the usual ++-polarization in GR. This clarifies that the GR solution exists as a particular case of the general GHMPG solution.
@@ III.3 Scalar modes
where $\Phi^{\pm}$ and $\Psi^{\pm}$ represent the amplitudes of the outgoing (++) and ingoing (-) waves, and $p^{\mu}$ and $q^{\mu}$ are the wave vectors. Due to the fact that the scalar modes are massive, the normalization conditions for the wave vectors in this case become
@@
$M_{\Phi}^{2}=\frac{1}{3}\left(A-\sqrt{B}\right),$ (42)
$M_{\Psi}^{2}=\frac{1}{3}\left(A+\sqrt{B}\right),$
@@ IV.1 The Newman-Penrose formalism
[…]
@@
Eardley, Lee, and Lightman have developed a formalism for classifying the gravitational wave solution for metric theories allowing for plane null gravitational wave solutions, that is, with massless modes, in which the NP quantities reduce to only six real independent components, namely the set $\{\Psi_{2},~\Psi_{3},~\Psi_{4},~\Phi_{22}\}$ (the other nonvanishing components can be shown to be linear combinations of the elements of this set). The associated symmetry group is a subgroup of the Lorentz group known as the “little group” $E(2)$ that leaves the wave-vector invariant. Under the action of the rotation group, the values of the helicity for each of the components of the set are, respectively, $\{0,~\pm 1,\pm 2,0\}$. Hence, $\Psi_{2}$ represents a longitudinal mode, $\Psi_{3}$ can be split into the real and purely imaginary parts responsible for the vectorial $y$- and $z$- modes, $\Psi_{4}$ can also be split as the former leading to the tensor ++ and $\times$ polarizations, and $\Phi_{22}$ represents the scalar breathing mode ELL1 ; ELL2 .
For gravity theories with massive gravitational wave polarization modes, the formalism outlined above cannot be straightforwardly applied. As we have found in Sec. III through the linearization method, the GHMPG features two massive scalar modes that can be decoupled. Thus, we shall use the method of inverting the NP quantities as functions of the polarization states as done in Ref. Polarisations .
@@ IV.2 Massive Propagating Modes
where $p_{1},\dots,p_{6}$ are the longitudinal, vectorial y-, vectorial z-, ++, $\times$, and breathing polarization modes, respectively.
@@
Equations (59) represent the most general description of gravitational wave polarization modes in the GHMPG, since the assumption of propagation in the $x$ direction is arbitrary. To proceed with the analysis of these modes, it is necessary to break generality either imposing explicit forms for the quantities appearing in the equations, e.g. choosing a particular form of the theory, or to impose additional constraints on these quantities through a comparison with observational data.
@@ V Constraints and implications
For the particular case of plane gravitational waves propagating along the $x$ direction, which we consider to be an acceptable approximation given the large distance to the sources of the gravitational wave signals detected, the displacement is constant along each plane orthogonal to the propagation direction. Hence we expect that $g_{xy}=0=g_{xz}$, a result that is in agreement with the linearized case (see Eq. (32)), and therefore no vector polarization modes arise in the full non-linear regime of the theory, i.e. $p_{2}^{(y)}=p_{3}^{(z)}=0$.
Additional simplifications to the results presented in Eqs. (59) can be obtained by taking into account the constrains on the scalar fields obtained from a cosmological analysis of the theory. Indeed, this gravity model has been thoroughly studied in a cosmological context through the dynamical systems approach Rosa:2024pzo , and recent results show that the fields $\varphi$ and $\psi$ are exponentially suppressed during the matter-dominated epoch, reaching approximately constant values at late times, i.e. $\varphi,\psi\approx\text{const}$. Furthermore, the solar system dynamics constrain the scalar fields to $\varphi-\psi\approx 1$. As such, the potential $V$ is responsible for late-time cosmic acceleration and plays the role of a cosmological constant at late times, whose value is constrained to $V\approx 10^{-122}/m_{P}^{2}$, with $m_{P}$ the Planck mass $m_{P}=\sqrt{\hbar G/c^{3}}$.
Taking the cosmological results into account at non-linear level, we expect that at present times, since $\partial_{\mu}\varphi\approx 0$, $\partial_{\mu}\psi\approx 0$, and $V\sim 10^{-122}$, and thus $\Xi\approx 0$, the theory may exhibit highly-suppressed scalar modes, $p_{1}^{(l)}$, $p_{6}^{(b)}$, and tensor polarizations which are essentially dominated by the Riemann tensor components. Even so, it is remarkable that the ++ polarization is not necessarily identical to the one in GR, as it is not required that $R_{xyxy}=-R_{xzxz}$, and hence we there is a possible ”smoking gun” for distinguishing between GHMPG and GR. However, it is important to note that since GR exists as a particular solution of GHMPG, in the absence of potentially measurable differences between $R_{xyxy}$ and $R_{xzxz}$, this analysis renders the GHMPG unfalsifiable from the point of view of gravitational wave propagation in the plane gravitational wave approximation.
@@
$p_{4}^{(+)}$ $=\pm\frac{1}{2}\omega^{2}\left(\delta g_{yy}-\delta g_{zz}\right)~,$ (62)
$p_{5}^{(\times)}$ $=\pm\frac{1}{2}\omega^{2}\delta g_{yz}~,$
$p_{6}^{(b)}$ $=-\frac{1}{2}\left(\delta g_{yy,tt}+\delta g_{zz,tt}\right)~.$
In fact, one could fully estimate the longitudinal mode in the linear limit by using Eqs. (38) and (40) to obtain